A person whose mass is m = 59.5 kg steps on a mechanical bathroom scale placed on an inclined plane that makes the angle ? = 25.5� with the horizontal. What is the reading on the scale?
Answer #1
Concepts and reason

The concepts used in solving this problem are resolution of vectors. The reading

on the bathroom scale is calculated by resolving the vectors as the scale is placed on the inclined plane.

Fundamentals

Resolving a vector, A is finding the components in mutually perpendicular directions. The magnitude of the two components can be written as AcosθA\cos \theta and AsinθA\sin \theta .

Asin 0
bo
Acos e

The pictorial description for the problem is as follows:

mg sino
A
mg cos

In the above figure, θ\theta is the angle made by the scale with the horizontal , m is the mass, g is the acceleration due to gravity, mgcosθmg\cos \theta is the horizontal component, and mgsinθmg\sin \theta is the vertical component of the scale

As the bathroom scale is placed on the inclined plane angled 25.525.5^\circ with the horizontal, the reading on the scale is given by,

W=mgcosθW = mg\cos \theta

Substitute 59.5 kg for mg and 25.525.5^\circ for θ\theta .

W=59.5kg(cos25.5)=53.70kg\begin{array}{c}\\W = 59.5\,{\rm{kg}}\left( {\cos 25.5^\circ } \right)\\\\ = 53.70{\rm{ kg}}\\\end{array}

Therefore, the reading on the bathroom scale is 53.70kg53.70{\rm{ kg}} .

Ans:

The reading on the bathroom scale is 53.70 kg.